### v1

# Evaluating Conversational Recommender Systems with Large Language Models

Abstract.
Conversational recommender systems (CRS) involve both recommendation and dialogue tasks, which makes their evaluation a unique challenge. Although past research has analyzed various factors that may affect user satisfaction with CRS interactions from the perspective of user studies, few evaluation metrics for CRS have been proposed. Recent studies have shown that LLMs can align with human preferences, and several LLM-based text quality evaluation measures have been introduced. However, the application of LLMs in CRS evaluation remains relatively limited. To address this research gap and advance the development of user-centric conversational recommender systems, this study proposes an automated LLM-based CRS evaluation framework, building upon existing research in human-computer interaction and psychology. The framework evaluates CRS from four dimensions: dialogue behavior, language expression, recommendation items, and response content. We use this framework to evaluate four different conversational recommender systems.
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Figure 1. Our proposed evaluation framework encompasses four dimensions, covering a total of 12 metrics. For a conversational recommendation system (CRS) and its interaction history with users, we employ large language models (LLMs) to perform automated scoring, resulting in scores for the 12 metrics.
[…]
Our proposed evaluation framework encompasses four dimensions, covering a total of 12 metrics. For a conversational recommendation system (CRS) and its interaction history with users, we employ large language models (LLMs) to perform automated scoring, resulting in scores for the 12 metrics.
In recent years, conversational recommendation have gained growing attention within the research community of recommender systems, and various implementations of Conversational Recommender Systems (CRSs) have been proposed (e.g.,  (Chen et al., 2019; Wang et al., 2022a; Sun and Zhang, 2018; Tu et al., 2022; Wang et al., 2022b)). Unlike traditional Recommender Systems (RS), which model user interests and generate recommendations based on implicit feedback signals such as clicking or browsing history  (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008; Bell and Koren, 2007; Koren et al., 2009; Covington et al., 2016), CRSs solicit and identify user interests through natural language conversations to make recommendations (Zamani et al., 2023; Jannach et al., 2021; Gao et al., 2021).
In conversational recommendation, a typical conversation between a user and a CRS is shown in Figure 1. The system generates text responses to the user and also provides an idependent list of recommended items. Different from traditional RS which are solely concerned with item recommendation, CRS not only address item recommendation but also manage conversations, i.e., engaging in multi-turn, semantically coherent dialogues.
The complexity of conversation recommendation tasks presents challenges for the evaluation of CRSs. Existing research often treats item recommendation and conversation management as two separate tasks when evaluating the performance of CRS, using corresponding metrics for each task (Chen et al., 2019; Wang et al., 2022a, b; Zhang et al., 2024; Feng et al., 2023). For example, Feng et al. (2023) evaluate CRS performance in item recommendation using Recall, and in conversation management using BLEU and distinct-n. However, such rule-based evaluation metrics for CRSs, such as Recall, BLEU and distinct-n, often do not align with actual user evaluations based on interactive experiences. Human-annotated evaluations help align CRSs with real user preferences; however, the cost of obtaining manually labeled data is extremely high.
In recent years, as large language models (LLMs) have demonstrated significant natural language understanding capabilities, researchers have proposed various methods to leverage LLMs for evaluating text quality and have demonstrated that LLMs have the potential to align with human preferences. (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023; Wang et al., 2024; Gao et al., 2023). However, to our knowledge, the use of LLMs for evaluating CRS remains quite limited.
To advance the development of user-centric conversational recommendation systems, in this work, we build upon existing literature and propose a framework of evaluation metrics for CRSs using LLMs from a user-centric perspective, considering four dimensions: (1) Dialogue Behavior, which primarily focuses on whether the CRS’s actions during interactions with users meet user expectations, whether it can correct errors based on user requests, and whether it can proactively inquire about user preferences. (2) Language Expression, which primarily focuses on whether the text generated by the CRS conforms to natural human language usage, whether there are any obvious grammatical errors, and whether the language used is appropriate. (3) Recommendation Items, which primarily focuses on whether the items recommended by the recommendation system meet the user’s needs and whether they also balance diversity and novelty based on this foundation. (4) Response Content, which primarily focuses on whether the text generated by the CRS during the dialogue is semantically closely related to the recommendation list provided by the CRS, whether it offers in-depth explanations of the recommendations within the text, and whether the content of the text is factually correct. Figure 1 presents the four evaluation dimensions we proposed and the corresponding metrics under each dimension.
We evaluated BARCOR (Wang et al., 2022a), CHATCRS (Wang et al., 2023a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) across the 12 factors in the proposed framework, using three LLMs: GPT-4o-mini-0718, LLaMa-3-8B, and GLM-4-Air, on the OpenDialKG (Moon et al., 2019) and ReDial (Li et al., 2019) datasets. Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets.
To our knowledge, we believe we are the first to propose a user-centric LLM-based evaluating framework for CRSs, incorporating insights from previous work in the field of recommender system, psychology, human-computer interaction. The main contributions of our work is that we systematically compiles and summarizes a comprehensive and diverse set of metrics for evaluating CRSs from a user-centric perspective, utilizing LLM-based standardized assessments. By using LLMs as agents for real users in multi-facet evaluations, our framework can consider various factors that influence user satisfaction and score CRS performance from different perspectives, bridging the gap between the fields of human-computer interaction and recommender system technology.
2. Related Work
2.1. Conversational Recommender Systems
Conversational recommender systems (CRSs) interact with users through multiple turns of dialogue, solicit and identify user interests through natural language conversations (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023). Most of the implementation methods are based on neural networks and combine techniques such as knowledge graphs (Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022a, b). In recent years, methods that directly utilize Large Language Models (LLMs) as CRS have also emerged (Wang et al., 2023a).
In these implementations, the interaction patterns between users and CRSs can be divided into two main categories: One line of work relies on handcrafted response templates and predefined interactive actions (e.g., asking users’ preferences for item attributes or making recommendations) to interact with users (e.g., (Lei et al., 2020; Tu et al., 2022; Zhou et al., 2020; Shang et al., 2023)); Another line of work focuses on interacting with users through more flexible natural language conversations, also known as chit-chat (e.g., (Chen et al., 2019; Wang et al., 2022b, a)). In our study, we primarily focus on the evaluation of chit-chat CRS.
We selected four representative CRSs for evaluation, The three methods are based on neural networks, namely BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b), while one directly utilizes GPT as a CRS (CHATCRS) (Wang et al., 2023a). We selected these systems because many studies ( e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)) have used them as baselines or evaluated them, which helps to connect our work with existing evaluations of conversational recommender systems.
2.2. Evaluating Conversational Recommender Systems
In fields such as Human-Computer Interaction (HCI) and psychology, there have been many research based on user studies analyzing what factors can influence users’ experience with CRS (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020). Nevertheless, only a few work have proposed metrics for evaluating CRSs (Yang et al., 2024; Huang et al., 2024).
In the RS community, most work still rely on metrics such as Recall and nDCG to measure the effectiveness of CRS recommendations, and metrics like BLEU and distinct-n to measure the quality of text generated by CRS.
Since obtaining interaction logs and feedback from real users is costly, a common method for evaluating Chit-chat CRS is user simulation (Wang et al., 2023a; Zhang and Balog, 2020). Wang et al. (Wang et al., 2023a) proposed a user simulator based on LLM and argued that it can be a reliable alternative to the human evaluator. In this work, we also adopt the user simulation-based approach for evaluation, using LLMs as user simulators to engage in multi-turn dialogues with the CRS, following the work of Wang et al. (Wang et al., 2023a).
3. CoRE: Our Proposed Method
Figure 2. The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 factors.
[…]
The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 dimensions.
Inspired by previous studies (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020), we summarize and list 12 factors that affect user satisfaction with interactive recommender systems, and categorize them into four dimensions based on the elements involved in user-system interactions.
Dialogue Behavior. This dimension focuses on whether the system can take actions that align with the user’s intentions during the conversation, in order to evaluate the performance of the CRS as a conversational agent. The main evaluation target is the system’s behavior in the dialogue. The factors under this dimension are also used to evaluate general conversational agents. The factors under this dimension include:
  • Coherence. Coherence refers to the system’s ability to understand the user’s intention and take action correspondingly, ensuring a logical and relevant flow of conversation. For example, when the user requests a recommendation, the system provides one; when the user asks the system to introduce a specific item, the system is able to do so.
  • Recoverability. Recoverability refers to the system’s ability to recognize and correct mistakes based on user feedback, adjusting its responses accordingly. This involves the system effectively understanding user feedback or corrections and adjusting its responses accordingly in the subsequent conversation.
  • Proactiveness. Proactiveness refers to the system’s ability to actively shape and guide the conversation. This involves the system not only responding to the user’s queries but also initiating topics, asking about user preferences, and suggesting relevant information or follow-up questions.
Language Expression.This dimension focuses on whether the system’s response text is sufficiently close to human expression, being fluent, natural, and appropriate in language. The main evaluation target is the text generated by the system. The factors in this dimension are also used to evaluate natural language generation tasks. The factors under this dimension include:
  • Grammatical correctness. Grammatical correctness refers to the degree to which the text adheres to standard grammatical rules, including proper sentence structure, subject-verb agreement, tense consistency, and correct usage of words and phrases.
  • Naturalness. Naturalness refers to how closely the system-generated text resembles native speaker expressions in vocabulary and grammar, ensuring it is easy to understand, avoiding overly complex or unusual words and grammatical constructions.
  • Appropriateness. Appropriateness refers to the system’s sensitivity to cultural, ethical, and social norms. The language used should be polite, respectful, and free from offensive or insulting content.
Recommendation Items. This dimension focuses on whether the items in the system’s recommended list meet the user’s needs while maintaining novelty and diversity. The main evaluation target is the recommendation list provided by the system. The factors of this dimension are also used to evaluate traditional RS. The factors under this dimension include:
  • Effectiveness. Effectiveness refers to how well the system’s recommended items align with the user’s expressed interests in a conversation.
  • Novelty. Novelty refers to the degree of freshness of an item. The higher the novelty, the less familiar the recommended item is to the user.
  • Diversity. Diversity in recommendations refers to the variety of items presented by the system, ensuring that the recommended list includes a range of different, yet relevant options. It is measured by the presence of different features of items in the recommendations.
Response Content. This dimension focuses on the CRS’s ability to integrate text generation with recommendation tasks. The main evaluation targets involve both the system-generated text and the recommended items. The factors in this dimension are specific to conversational recommendation tasks, distinguishing them from traditional text generation, dialogue agent, or recommendation tasks. The factors under this dimension include:
  • Semantic Relevance. Semantic relevance refers to the degree of connection between the content generated by the system and the items mentioned in the recommendation list provided by the system, or to what extent is the content discussed by the system closely related to the items in the recommendation list.
  • Explainability.Explainability refers to the system’s ability to provide clear and detailed explanations for its recommendations. The explanation should be persuasive, helping the user understand why certain items are suggested and how well these items match their preferences.
  • Groundness. Groundness refers to the factual accuracy of system-generated text, ensuring that the descriptions of items are correct, based on verifiable facts, and free from false or misleading information.
We use LLMs as evaluators, providing scores from 0 to 4 based on the 12 factors mentioned above. Figure 2 illustrates the overall process of our method; Figure 1 shows how the 12 factors correspond to elements in the user-CRS interaction.
4. Experiments
Dataset #Dialogues #Utterances Domains
ReDial 10,006 182,150 Movie
OpenDialKG 13,802 91,209 Movie, Book, Sports, Music
Table 1. Summary of CRS Datasets
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.66 (0.66) 3.79 (0.61) 4.0 (0.0) 4.0 (0.04) 2.4 (1.38) 3.39 (0.96) 3.25 (1.13) 3.7 (0.72)
Coherence 1.19 (0.72) 1.34 (0.77) 3.74 (0.5) 3.72 (0.66) 0.47 (0.62) 1.17 (0.68) 0.51 (0.59) 0.77 (0.66)
Diversity 3.5 (0.7) 3.38 (0.78) 3.15 (0.9) 3.03 (0.89) 3.24 (0.94) 3.13 (0.68) 3.14 (0.93) 3.4 (0.77)
Effectiveness 1.88 (0.65) 2.33 (0.93) 3.71 (0.7) 3.69 (0.71) 1.11 (0.32) 1.7 (0.61) 1.42 (0.52) 1.79 (0.62)
Explainability 0.5 (0.54) 0.36 (0.5) 2.8 (1.42) 2.52 (1.42) 0.02 (0.15) 0.1 (0.3) 0.16 (0.37) 0.26 (0.45)
Grammar 2.12 (0.98) 2.92 (1.0) 3.91 (0.31) 3.9 (0.36) 0.55 (0.76) 2.79 (1.14) 2.21 (1.08) 2.87 (0.97)
Groundness 1.25 (0.81) 3.13 (0.67) 3.68 (0.71) 3.82 (0.51) 0.55 (0.65) 1.75 (1.34) 0.88 (0.79) 1.97 (1.11)
Naturalness 2.29 (0.76) 3.13 (0.67) 3.93 (0.25) 3.93 (0.26) 0.94 (0.5) 2.73 (0.93) 1.87 (0.73) 2.68 (0.71)
Novelty 1.2 (0.95) 0.25 (0.52) 1.55 (1.06) 1.12 (0.97) 1.47 (1.08) 0.76 (0.97) 1.37 (1.04) 0.76 (0.75)
Proactiveness 0.96 (0.6) 1.07 (0.63) 3.64 (0.52) 3.67 (0.53) 0.12 (0.86) 0.52 (0.54) 0.42 (0.51) 0.59 (0.53)
Recoverability 2.12 (1.33) 2.05 (1.53) 3.82 (0.52) 3.8 (0.63) 1.06 (1.06) 2.08 (1.16) 1.05 (1.07) 1.37 (1.15)
Semantic Rel. 1.11 (1.52) 1.37 (1.53) 2.79 (1.47) 2.72 (1.3) 0.67 (1.06) 0.59 (1.17) 1.46 (1.46) 1.82 (1.56)
Table 2. Evaluation result given by GPT-4o-mini-0718 in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.22 (1.43) 3.77 (0.87) 4.00 (0.12) 3.99 (0.11) 0.64 (1.05) 3.18 (1.47) 1.86 (1.81) 3.20 (1.48)
Coherence 1.50 (0.59) 1.64 (0.67) 3.78 (0.51) 3.77 (0.51) 1.05 (0.37) 1.26 (0.51) 1.15 (0.46) 1.25 (0.57)
Diversity 1.82 (0.92) 2.47 (0.97) 3.05 (0.99) 3.33 (0.82) 1.56 (0.93) 2.47 (1.00) 1.87 (0.84) 2.11 (0.88)
Effectiveness 1.98 (1.20) 2.12 (1.09) 3.66 (0.69) 3.71 (0.70) 1.15 (1.09) 1.70 (1.13) 1.69 (1.11) 1.85 (1.15)
Explainability 0.44 (0.57) 0.54 (0.76) 3.69 (0.85) 3.58 (1.03) 0.00 (0.06) 0.07 (0.27) 0.18 (0.39) 0.28 (0.47)
Grammar 2.20 (0.72) 2.78 (0.90) 3.87 (0.35) 3.86 (0.36) 1.14 (0.36) 2.31 (0.96) 1.84 (0.73) 2.37 (0.86)
Groundness 1.28 (0.62) 2.31 (1.24) 3.90 (0.41) 3.92 (0.36) 0.97 (0.27) 1.92 (1.37) 1.06 (0.39) 1.71 (1.12)
Naturalness 2.22 (0.87) 2.90 (0.99) 3.63 (0.54) 3.63 (0.55) 1.02 (0.27) 2.53 (1.07) 1.62 (0.79) 2.52 (0.92)
Novelty 2.01 (0.82) 2.25 (0.79) 2.53 (0.76) 2.66 (0.72) 1.85 (0.77) 2.40 (0.75) 2.00 (0.65) 2.01 (0.85)
Proactiveness 1.05 (0.60) 1.01 (0.64) 3.74 (0.56) 3.65 (0.70) 0.31 (0.46) 0.47 (0.54) 0.72 (0.53) 0.71 (0.52)
Recoverability 1.87 (0.53) 2.16 (0.80) 3.85 (0.50) 3.85 (0.49) 1.26 (0.46) 1.93 (0.69) 1.41 (0.53) 1.67 (0.73)
Semantic Rel. 0.96 (1.08) 1.45 (1.06) 2.68 (1.22) 2.83 (1.11) 0.47 (0.76) 0.83 (1.04) 1.27 (1.06) 1.40 (1.23)
Table 3. Evaluation result given by LLaMa-3-8B in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.83 (0.53) 3.95 (0.28) 4.00 (0.00) 4.00 (0.04) 2.50 (1.05) 3.46 (0.88) 3.30 (0.99) 3.59 (0.78)
Coherence 1.44 (1.04) 1.55 (1.09) 3.88 (0.41) 3.79 (0.57) 0.53 (0.81) 1.35 (0.99) 0.48 (0.83) 0.69 (0.98)
Diversity 2.63 (1.29) 3.34 (1.07) 3.25 (0.91) 3.48 (0.75) 2.63 (1.34) 3.30 (1.02) 2.83 (1.20) 3.05 (1.37)
Effectiveness 1.87 (1.49) 2.03 (1.29) 3.69 (0.66) 3.74 (0.60) 1.24 (1.36) 1.59 (1.33) 1.36 (1.34) 2.05 (1.38)
Explainability 0.86 (0.67) 0.64 (0.67) 2.88 (1.04) 2.70 (1.15) 0.50 (0.38) 0.36 (0.50) 0.50 (0.58) 0.49 (0.59)
Grammar 2.52 (0.89) 3.24 (0.77) 3.85 (0.37) 3.81 (0.42) 1.28 (0.99) 2.88 (1.07) 2.48 (1.03) 3.06 (0.83)
Groundness 1.48 (1.02) 2.71 (1.00) 3.74 (0.43) 3.87 (0.47) 1.08 (1.00) 2.69 (1.12) 1.04 (1.04) 2.57 (1.08)
Naturalness 1.68 (0.70) 3.27 (0.57) 3.76 (0.72) 3.74 (0.44) 1.76 (0.65) 2.90 (0.86) 2.29 (0.71) 2.81 (0.63)
Novelty 2.26 (1.25) 1.33 (1.19) 3.01 (1.14) 2.64 (1.22) 2.02 (1.42) 1.71 (1.26) 2.57 (1.43) 1.42 (1.30)
Proactiveness 1.12 (0.73) 1.26 (0.76) 3.47 (0.61) 3.43 (0.62) 0.51 (0.52) 0.81 (0.52) 0.69 (0.58) 0.69 (0.58)
Recoverability 2.33 (1.32) 2.22 (1.36) 3.91 (0.46) 3.85 (0.55) 1.22 (1.15) 2.74 (0.96) 1.11 (1.14) 1.20 (1.39)
Semantic Rel. 1.33 (1.20) 1.56 (1.28) 2.39 (1.32) 2.90 (1.06) 0.52 (0.88) 0.73 (1.07) 1.33 (1.22) 1.53 (1.36)
Table 4. Evaluation result given by GLM-4-Air in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
4.1. Experimental Setup
Dataset. Following previous work (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), we select ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019) as the benchmark datasets. ReDial primarily focuses on movie recommendations, and OpenDialKG encompasses a broader range of domains, including movies, books, sports, and music. Table 1 shows a summary of the two datasets.
CRS Model. Following previous work (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), we present a comparative evaluation and analysis towards four CRS: BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) and CHATCRS (Wang et al., 2023a). For BARCOR, KBRD, and UniCRS, we use the pre-trained models provided by Wang et al (Wang et al., 2023a); For CHATCRS, we implemented several modifications, utilizing the GPT-4o-mini-0718 as the LLM model.
User Simulator. Building on the work of Wang et al. (Wang et al., 2023a), we use GPT-4o-mini-0718 to simulate users interacting with CRS in dialogues. To ensure sufficient system dialogue text, we modified the interaction rules and prompts to ensure that each dialogue session between the user simulator and CRS includes at least three and no more than five rounds of interaction. More specifically, before each dialogue session begins, we input the ground truth item from a real user into the simulator, requiring the simulator to describe relevant attributes in the dialogue without directly revealing the ground truth item. In each round of dialogue, the system generates response text and a list of recommended items. If the system identifies the ground-truth item, and more than three rounds of dialogue have occurred, the session ends; if more than five rounds of dialogue have occurred and the system has not identified the ground-truth item in any of the previous rounds, the session also ends.
Evaluation Process. After obtaining dialogue logs through user simulation, we post-processed the logs by consolidating the items recommended by CRS in each round into a session-level set (for the sake of convention, we refer to the set as the “recommendation list” in the paper). The left side of Figure 2 shows an example of a processed log, where all items recommended by the system during the dialogue are combined into the recommendation list. For each CRS, we obtained 3,341 records on the ReDial evaluation set and 1,187 records on the OpenDialKG evaluation set.
We then utilized LLMs to evaluate the dialogue sessions in the logs, based on the 12 factors listed in Section 3. For each factor, we have described its definition, scoring criteria, and scoring steps in the instruction. For each factor, a score between 0 and 4 was obtained. For details on the scoring, refer to Appendix A. We selected three LLMs as Evaluators, namely GPT-4o-mini-0718, LLaMa3-Chat-8B, and GLM-4-Air.
4.2. Experimental Result
Table 23, and  4 shows the evaluation results of four systems on two datasets are presented separately, using GPT-4o-mini-0718, LLaMa3-8B, and GLM-4-Air as evaluators.
From Tables 2,3, and4, we can reach the following findings regarding the rankings of systems:
  1. (1)
    Across both the ReDial and OpenDialKG datasets, the three Evaluators consistently give the highest scores to CHATCRS on most metrics. Both LLaMa and GLM give CHATCRS the highest scores on all metrics, while GPT gives CHATCRS the highest scores on all factors except for Diversity.
  2. (2)
    Across both the ReDial and OpenDialKG datasets, the three Evaluators consistently give BARCOR the second-highest scores on most metrics. LLaMa gives BARCOR the second-highest score on all metrics except for Novelty on ReDial; on OpenDialKG, it gives BARCOR the second-highest score on all metrics except for Diversity and Semantic Relevance. According to the scores from GPT and GLM, BARCOR ranks second in the scores for at least eight factors across both datasets.
  3. (3)
    For the systems ranked third and fourth in each factor, there is a significant amount of inconsistency in the results given by different evaluators across different datasets. For example, in most cases, both GPT and LLaMa gave UniCRS the third-highest score and KBRD the lowest score. However, this pattern does not apply to GLM. GLM gave KBRD the third-highest score in Coherence and Groundness, while giving UniCRS the lowest score. In Novelty, GLM ranked the four systems on the OpenDialKG dataset as CHATCRS ¿ UniCRS ¿ BARCOR ¿ KBRD, but on the ReDial dataset, the ranking was CHATCRS ¿ KBRD ¿ UniCRS ¿ BARCOR.
Further observation of the score rankings given by the three evaluators for the four systems, we reach the following findings reagarding the consistency (or inconsistency) of the system score rankings among the three evaluators:
  1. (1)
    In most cases, the score rankings given by GPT and LLaMa are consistent. The system score rankings for Appropriateness, Coherence, Effectiveness, Explainability, Naturalness, and Proactiveness given by GPT and LLaMa are consistent across the two datasets. For the factors of Grammar, Groundness, and Semantic Relevance, GPT and LLaMa also achieve consistent rankings for the systems on one dataset.
  2. (2)
    The score rankings given by GLM are not as consistent with those of GPT or LLaMa. GLM is only consistent with GPT and LLaMa on both datasets in terms of the score rankings for the Appropriateness factor on the OpenDialKG dataset. For the factors of Coherence, Effectiveness, Explainability, Naturalness, and Proactiveness, where GPT and LLaMa have reached consensus on the rankings across both datasets, GLM can only achieve the same conclusions as GPT or LLaMa on the ReDial dataset.
From the perspective of the datasets, we found that GPT and LLaMa tend to give more consistent rankings on OpenDialKG, while GPT and GLM tend to give more consistent rankings on ReDial.
Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets. We also found that: In most cases, the score rankings given by GPT and LLaMa are consistent; but the score rankings given by GLM are not as consistent with those of GPT or LLaMa.
5. Conclusion
5.1. Main Findings
Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets. We also found that: In most cases, the score rankings given by GPT and LLaMa are consistent; but the score rankings given by GLM are not as consistent with those of GPT or LLaMa.
We believe there are two possible reasons for the inconsistency in LLMs’ scores. First, each model learns different parameterized knowledge during training due to variations in the training corpora, which can affect their judgment. Second, some factors, such as diversity or novelty, are more subjective in their criteria compared to others, such as groundness or grammar, which can also influence the scores given by LLMs, reflected in the system rankings and the standard deviation of each system’s score.
5.2. Limitations and Future Directions
Firstly, our work has not examined how well LLM evaluators’ scores align with human ratings, and future research should include user studies and human feedback. Secondly, we have not yet addressed the gap in CRS evaluation where recommendation and conversational tasks are treated separately and need to be integrated into a unified approach.
Additionally, previous work on search system evaluation has shown that users may be influenced by cognitive biases during their interactions with systems (Chen et al., 2023a, 2024). Future research needs to further integrate psychological theories to make the evaluation results closer to real user feedback.
Appendix A Evaluation Instructions
A.1. Instruction Template
Figure 3. The instruction templates we provide to the large models.
Figure 3 shows the instruction templates we provided to the LLM evaluators. The content of FACTOR_DEFINITION, EVALUATION_STANDARD, EVALUATION_STEP of each factor and an example of CONVERSATION_LOG can be found below.
A.2. Evaluation Guidelines
Figure 4567,89101112131415 shows the evaluation guidelines of each factor we provided to the LLM evaluators respectively.
Figure 4. Evaluation guideline of appropriateness
Figure 5. Evaluation guideline of coherence
Figure 6. Evaluation guideline of diversity
Figure 7. Evaluation guideline of effectiveness
Figure 8. Evaluation guideline of explainability
Figure 9. Evaluation guideline of grammatical correctnes
Figure 10. Evaluation guideline of groundness
Figure 11. Evaluation guideline of naturalness
Figure 12. Evaluation guideline of novelty
Figure 13. Evaluation guideline of proactiveness
Figure 14. Evaluation guideline of recoverability
Figure 15. Evaluation guideline of semantic relevance
Figure 16. An example of the conversation log after processing
A.3. Example of Conversation Log
Figure 16 shows an example of the conversation log after processing.