### v2

# Large Language Models as Evaluators for Conversational Recommender Systems

Abstract.
Conversational recommender systems (CRS) involve both recommendation and dialogue tasks, which makes their evaluation a unique challenge. Although past research has analyzed various factors that may affect user satisfaction with CRS interactions from the perspective of user studies, few evaluation metrics for CRS have been proposed. Recent studies have shown that LLMs can align with human preferences, and several LLM-based text quality evaluation measures have been introduced. However, the application of LLMs in CRS evaluation remains relatively limited. To address this research gap and advance the development of user-centric conversational recommender systems, this study proposes an automated LLM-based CRS evaluation framework, building upon existing research in human-computer interaction and psychology. The framework evaluates CRS from four dimensions: dialogue behavior, language expression, recommendation items, and response content. We use this framework to evaluate four different conversational recommender systems.
Conversational Recommender Systems, Information Access, Dialogue Systems,Information System Evaluation
copyright: noneconference: The 1st Workshop on Human-Centered Recommender Systems (HCRS) at TheWebConf 2025; 28 April - 2 May, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
In recent years, conversational recommendation have gained growing attention within the research community of recommender systems, and various implementations of Conversational Recommender Systems (CRSs) have been proposed (Chen et al., 2019; Wang et al., 2022a; Sun and Zhang, 2018; Tu et al., 2022; Wang et al., 2022b). Unlike traditional Recommender Systems (RS), which model user interests and generate recommendations based on implicit feedback signals such as clicking or browsing history  (Koren, 2008; Mao et al., 2021; Zhu et al., 2023, 2022), CRSs solicit and identify user interests through natural language conversations to make recommendations (Zamani et al., 2023; Jannach et al., 2021; Gao et al., 2021).
Figure 1. Our proposed evaluation framework encompasses four dimensions, covering a total of 12 metrics. For a conversational recommendation system (CRS) and its interaction history with users, we employ large language models (LLMs) to perform automated scoring, resulting in scores for the 12 metrics.
[…]
Our proposed evaluation framework encompasses four dimensions, covering a total of 12 metrics. For a conversational recommendation system (CRS) and its interaction history with users, we employ large language models (LLMs) to perform automated scoring, resulting in scores for the 12 metrics.
In conversational recommendation, a typical conversation between a user and a CRS is shown in Figure 1. The system generates text responses to the user and also provides an idependent list of recommended items. Different from traditional RS that focus solely on item recommendation, CRS not only provide item recommendation but also manage conversations, i.e., engaging in multi-turn, semantically coherent dialogues.
The complexity of conversation recommendation tasks presents challenges for the evaluation of CRSs. Existing research often treats item recommendation and conversation management as two separate tasks when evaluating the performance of CRS, using corresponding metrics for each task (Chen et al., 2019; Wang et al., 2022a, b; Zhang et al., 2024; Feng et al., 2023). For example, Feng et al. (2023) evaluate CRS performance in item recommendation using Recall, and in conversation management using BLEU and distinct-n. However, such rule-based evaluation metrics for CRSs, such as Recall, BLEU and distinct-n, often do not align with actual user evaluations based on interactive experiences. Human-annotated evaluations help align CRSs with real user preferences; however, the cost of obtaining manually labeled data is extremely high.
In recent years, large language models (LLMs) have shown significant capabilities in natural language understanding. As a result, researchers have proposed various methods for leveraging LLMs to evaluate text quality, demonstrating their potential to align closely with human preferences (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023; Wang et al., 2024; Gao et al., 2023). However, to the best our knowledge, the application of LLMs for evaluating CRS remains relatively underexplored.
To advance the development of conversational recommendation systems, this work builds upon existing literature and proposes a user-centric evaluation framework based on LLMs, which can comprehensively and naturally quantify users’ satisfaction with the interacted CRS. Specifically, within the framework, we first generate interaction dialogues between users and the system by substituting real humans with user simulators to avoid high cost as (Wang et al., 2023a). We then evaluate the proposed metrics based on the logs by prompting the LLM.
Our evaluation framework consists of 12 evaluation metrics of four dimensions, including dialogue behavior, language expression, recommendation items, and response content, as shown in Figure 1. We describe these four dimensions as follows:
  • Dialogue Behavior. This dimension focuses on whether the CRS’s actions during interactions align with user expectations, its ability to correct errors based on user requests, and whether it proactively inquires about user preferences.
  • Language Expression. This dimension evaluates whether the text generated by the CRS follows natural human language usage, is free of obvious grammatical errors, and whether the language is appropriate in context.
  • Recommendation Items. This dimension assesses whether the items recommended by the CRS meet the user’s preferences and needs while balancing diversity and novelty.
  • Response Content. This dimension examines whether the text generated by the CRS during the dialogue is semantically relevant to the recommendation list, whether it provides in-depth explanations of the recommendations, and whether the content is factually accurate.
We conducted evaluations of BARCOR (Wang et al., 2022a), CHATCRS (Wang et al., 2023a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) according to the 12 metrics within the proposed framework. The evaluations were carried out using three LLMs, namely GPT-4o-mini-0718, LLaMa-3-8B, and GLM-4-Air, on the OpenDialKG (Moon et al., 2019) and ReDial (Li et al., 2019) datasets. Overall, our findings indicate that: generally, the three evaluators assign the highest score to CHATCRS, followed by BARCOR with the second highest score. However, the rankings of KBRD and UniCRS fluctuate based on the specific factors and datasets under consideration.
To the best of our knowledge, we are the first to propose a user-centric evaluation framework for CRSs based on LLMs. This framework incorporates valuable insights from previous research in the fields of recommender systems, psychology, and human-computer interaction. The main contributions of our work are as follows: First, we have systematically compiled and summarized a comprehensive and diverse set of metrics for evaluating CRSs from a user-centric perspective. These metrics are evaluated through LLM-based standardized assessments. Second, by employing LLM-based user simualtors as surrogates for real users in multi-faceted evaluations, our framework can take into account a wide range of factors that influence user satisfaction in a highly efficient manner. It is thus able to score the performance of CRSs from different viewpoints, effectively bridging the gap between the fields of human-computer interaction and recommender systems.
2. Related Work
Figure 2. The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 factors.
[…]
The overview of our method. We evaluate user interactions with conversational systems through a large language model across 12 dimensions.
2.1. Conversational Recommender Systems
Conversational recommender systems (CRSs) interact with users through multiple turns of dialogue, soliciting and identifying user interests through natural language conversations (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023). Most implementation methods are based on neural networks and incorporate techniques such as knowledge graphs (Chen et al., 2019; Zhou et al., 2020b; Wang et al., 2022a, b). In recent years, methods that directly utilize Large Language Models as CRSs have also emerged (Wang et al., 2023a).
In these implementations, the interaction patterns between users and CRSs can be classified into two main categories. One approach relies on handcrafted response templates and predefined interactive actions (e.g., asking users about item attributes or making recommendations) to engage with users (Lei et al., 2020; Tu et al., 2022; Zhou et al., 2020b; Shang et al., 2023). The other approach focuses on more flexible interactions through natural language conversations, commonly referred to as chit-chat (Chen et al., 2019; Wang et al., 2022b, a). In our study, we primarily focus on evaluating chit-chat CRSs.
We selected four representative CRSs for evaluation. Three of these methods are based on neural networks, i.e., BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b), while one directly utilizes GPT as a CRS, namely CHATCRS (Wang et al., 2023a). These systems were chosen because they have been widely used as baselines or evaluated in previous studies (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), which allows our work to build upon and connect with existing evaluations of conversational recommender systems.
2.2. Evaluating Conversational Recommender Systems
In fields such as Human-Computer Interaction (HCI) and psychology, numerous studies have analyzed the factors influencing users’ experiences with CRSs (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020). However, few studies have formally proposed metrics for evaluating CRSs (Yang et al., 2024; Huang et al., 2024).
In the community of recommender systems, traditional evaluation methods for CRSs typically focus on assessing recommendation accuracy (with metrics such as Recall) and nDCG and dialogue response quality (with metrics such as BLEU and distinct-n) in isolation (Wang et al., 2022a; Zhou et al., 2020a). This approach significantly overlooks the role of the dynamic interaction process in shaping user experiences, thereby failing to accurately capture the true performance of a CRS (Wang et al., 2023a). Although online user testing is considered the gold standard for evaluating CRSs, due to its ability to capture real user interactions and feedback, its high time demands and substantial labor costs make it difficult to scale (Gao et al., 2021).
To tackle this problem, user simulators have emerged as a substitute of real humans to interact with CRSs for performance evaluation (Zhang and Balog, 2020; Afzali et al., 2023; Chen et al., 2025). Recently, LLM-based user simulators have been continuously proposed, including single-prompt methods (Wang et al., 2023a; Zhu et al., 2024b; Kim et al., 2024) and agent-based architectures (Zhu et al., 2024a; Chen et al., 2025). However, these existing simulator-based methods only use a very limited set of metrics, which is insufficient to comprehensively evaluate a CRS. In this work, we build upon these advancements by leveraging LLM-based user simulators in our evaluation framework, and further propose a comprehensive evaluation metrics to evaluate the interaction performance based on user-system interaction dialogues. We also propose to use LLMs as evluators for these metrics.
3. The Proposed Method
In this paper, we propose a user-centric evaluation framework based LLMs for conversational recommender systems as shown by Figure 2. It first employs LLM-based user simulators to interact with CRS under various contexts to obtain a diverse set of interaction dialogues, and then utilizes LLM as evaluator to score the carefully designed evaluation metrics (ranging from 0 to 4 in our experiments).
Inspired by previous studies (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020), we summarize and list 12 factors that affect user satisfaction with interactive recommender systems, and categorize them into four dimensions based on the elements involved in user-system interactions.
Dialogue Behavior. This dimension focuses on whether the system can take actions that align with the user’s intentions during the conversation, in order to evaluate the performance of the CRS as a conversational agent. The main evaluation target is the system’s behavior in the dialogue. The factors under this dimension are also used to evaluate general conversational agents. The factors under this dimension include:
  • Coherence. Coherence refers to the system’s ability to understand the user’s intention and take action correspondingly, ensuring a logical and relevant flow of conversation. For example, when the user requests a recommendation, the system provides one; when the user asks the system to introduce a specific item, the system is able to do so.
  • Recoverability. Recoverability refers to the system’s ability to recognize and correct mistakes based on user feedback, adjusting its responses accordingly. This involves the system effectively understanding user feedback or corrections and adjusting its responses accordingly in the subsequent conversation.
  • Proactiveness. Proactiveness refers to the system’s ability to actively shape and guide the conversation. This involves the system not only responding to the user’s queries but also initiating topics, asking about user preferences, and suggesting relevant information or follow-up questions.
Language Expression.This dimension focuses on whether the system’s response text is sufficiently close to human expression, being fluent, natural, and appropriate in language. The main evaluation target is the text generated by the system. The factors in this dimension are also used to evaluate natural language generation tasks. The factors under this dimension include:
  • Grammatical correctness. Grammatical correctness refers to the degree to which the text adheres to standard grammatical rules, including proper sentence structure, subject-verb agreement, tense consistency, and correct usage of words and phrases.
  • Naturalness. Naturalness refers to how closely the system-generated text resembles native speaker expressions in vocabulary and grammar, ensuring it is easy to understand, avoiding overly complex or unusual words and grammatical constructions.
  • Appropriateness. Appropriateness refers to the system’s sensitivity to cultural, ethical, and social norms. The language used should be polite, respectful, and free from offensive or insulting content.
Recommendation Items. This dimension focuses on whether the items in the system’s recommended list meet the user’s needs while maintaining novelty and diversity. The main evaluation target is the recommendation list provided by the system. The factors of this dimension are also used to evaluate traditional RS. The factors under this dimension include:
  • Effectiveness. Effectiveness refers to how well the system’s recommended items align with the user’s expressed interests in a conversation.
  • Novelty. Novelty refers to the degree of freshness of an item. The higher the novelty, the less familiar the recommended item is to the user.
  • Diversity. Diversity in recommendations refers to the variety of items presented by the system, ensuring that the recommended list includes a range of different, yet relevant options. It is measured by the presence of different features of items in the recommendations.
Response Content. This dimension focuses on the CRS’s ability to integrate text generation with recommendation tasks. The main evaluation targets involve both the system-generated text and the recommended items. The factors in this dimension are specific to conversational recommendation tasks, distinguishing them from traditional text generation, dialogue agent, or recommendation tasks. The factors under this dimension include:
  • Semantic Relevance. Semantic relevance refers to the degree of connection between the content generated by the system and the items mentioned in the recommendation list provided by the system, or to what extent is the content discussed by the system closely related to the items in the recommendation list.
  • Explainability.Explainability refers to the system’s ability to provide clear and detailed explanations for its recommendations. The explanation should be persuasive, helping the user understand why certain items are suggested and how well these items match their preferences.
  • Groundness. Groundness refers to the factual accuracy of system-generated text, ensuring that the descriptions of items are correct, based on verifiable facts, and free from false or misleading information.
We use LLMs as evaluators to assign scores from 0 to 4 for the 12 factors mentioned above separately, based on the carefully designed prompts as detailedly presented in Appendix A.
4. Experiments
Table 1. Summary of CRS Datasets
Dataset #Dialogues #Utterances Domains
ReDial 10,006 182,150 Movie
OpenDialKG 13,802 91,209 Movie, Book, Sports, Music
Table 2. Evaluation result given by GPT-4o-mini-0718 in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.66 (0.66) 3.79 (0.61) 4.0 (0.0) 4.0 (0.04) 2.4 (1.38) 3.39 (0.96) 3.25 (1.13) 3.7 (0.72)
Coherence 1.19 (0.72) 1.34 (0.77) 3.74 (0.5) 3.72 (0.66) 0.47 (0.62) 1.17 (0.68) 0.51 (0.59) 0.77 (0.66)
Diversity 3.5 (0.7) 3.38 (0.78) 3.15 (0.9) 3.03 (0.89) 3.24 (0.94) 3.13 (0.68) 3.14 (0.93) 3.4 (0.77)
Effectiveness 1.88 (0.65) 2.33 (0.93) 3.71 (0.7) 3.69 (0.71) 1.11 (0.32) 1.7 (0.61) 1.42 (0.52) 1.79 (0.62)
Explainability 0.5 (0.54) 0.36 (0.5) 2.8 (1.42) 2.52 (1.42) 0.02 (0.15) 0.1 (0.3) 0.16 (0.37) 0.26 (0.45)
Grammar 2.12 (0.98) 2.92 (1.0) 3.91 (0.31) 3.9 (0.36) 0.55 (0.76) 2.79 (1.14) 2.21 (1.08) 2.87 (0.97)
Groundness 1.25 (0.81) 3.13 (0.67) 3.68 (0.71) 3.82 (0.51) 0.55 (0.65) 1.75 (1.34) 0.88 (0.79) 1.97 (1.11)
Naturalness 2.29 (0.76) 3.13 (0.67) 3.93 (0.25) 3.93 (0.26) 0.94 (0.5) 2.73 (0.93) 1.87 (0.73) 2.68 (0.71)
Novelty 1.2 (0.95) 0.25 (0.52) 1.55 (1.06) 1.12 (0.97) 1.47 (1.08) 0.76 (0.97) 1.37 (1.04) 0.76 (0.75)
Proactiveness 0.96 (0.6) 1.07 (0.63) 3.64 (0.52) 3.67 (0.53) 0.12 (0.86) 0.52 (0.54) 0.42 (0.51) 0.59 (0.53)
Recoverability 2.12 (1.33) 2.05 (1.53) 3.82 (0.52) 3.8 (0.63) 1.06 (1.06) 2.08 (1.16) 1.05 (1.07) 1.37 (1.15)
Semantic Rel. 1.11 (1.52) 1.37 (1.53) 2.79 (1.47) 2.72 (1.3) 0.67 (1.06) 0.59 (1.17) 1.46 (1.46) 1.82 (1.56)
Table 3. Evaluation result given by LLaMa-3-8B in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.22 (1.43) 3.77 (0.87) 4.00 (0.12) 3.99 (0.11) 0.64 (1.05) 3.18 (1.47) 1.86 (1.81) 3.20 (1.48)
Coherence 1.50 (0.59) 1.64 (0.67) 3.78 (0.51) 3.77 (0.51) 1.05 (0.37) 1.26 (0.51) 1.15 (0.46) 1.25 (0.57)
Diversity 1.82 (0.92) 2.47 (0.97) 3.05 (0.99) 3.33 (0.82) 1.56 (0.93) 2.47 (1.00) 1.87 (0.84) 2.11 (0.88)
Effectiveness 1.98 (1.20) 2.12 (1.09) 3.66 (0.69) 3.71 (0.70) 1.15 (1.09) 1.70 (1.13) 1.69 (1.11) 1.85 (1.15)
Explainability 0.44 (0.57) 0.54 (0.76) 3.69 (0.85) 3.58 (1.03) 0.00 (0.06) 0.07 (0.27) 0.18 (0.39) 0.28 (0.47)
Grammar 2.20 (0.72) 2.78 (0.90) 3.87 (0.35) 3.86 (0.36) 1.14 (0.36) 2.31 (0.96) 1.84 (0.73) 2.37 (0.86)
Groundness 1.28 (0.62) 2.31 (1.24) 3.90 (0.41) 3.92 (0.36) 0.97 (0.27) 1.92 (1.37) 1.06 (0.39) 1.71 (1.12)
Naturalness 2.22 (0.87) 2.90 (0.99) 3.63 (0.54) 3.63 (0.55) 1.02 (0.27) 2.53 (1.07) 1.62 (0.79) 2.52 (0.92)
Novelty 2.01 (0.82) 2.25 (0.79) 2.53 (0.76) 2.66 (0.72) 1.85 (0.77) 2.40 (0.75) 2.00 (0.65) 2.01 (0.85)
Proactiveness 1.05 (0.60) 1.01 (0.64) 3.74 (0.56) 3.65 (0.70) 0.31 (0.46) 0.47 (0.54) 0.72 (0.53) 0.71 (0.52)
Recoverability 1.87 (0.53) 2.16 (0.80) 3.85 (0.50) 3.85 (0.49) 1.26 (0.46) 1.93 (0.69) 1.41 (0.53) 1.67 (0.73)
Semantic Rel. 0.96 (1.08) 1.45 (1.06) 2.68 (1.22) 2.83 (1.11) 0.47 (0.76) 0.83 (1.04) 1.27 (1.06) 1.40 (1.23)
Table 4. Evaluation result given by GLM-4-Air in the form of mean and standard deviation, with the number in parentheses representing the standard deviation. Bold, underline, and italics represent the highest, second-highest, and third-highest average scores of the systems on the corresponding dataset, respectively. If the average scores are the same, the system with the smaller standard deviation ranks higher.
System BARCOR CHATCRS KBRD UniCRS
Dataset OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial OpenDialKG ReDial
Appropriateness 3.83 (0.53) 3.95 (0.28) 4.00 (0.00) 4.00 (0.04) 2.50 (1.05) 3.46 (0.88) 3.30 (0.99) 3.59 (0.78)
Coherence 1.44 (1.04) 1.55 (1.09) 3.88 (0.41) 3.79 (0.57) 0.53 (0.81) 1.35 (0.99) 0.48 (0.83) 0.69 (0.98)
Diversity 2.63 (1.29) 3.34 (1.07) 3.25 (0.91) 3.48 (0.75) 2.63 (1.34) 3.30 (1.02) 2.83 (1.20) 3.05 (1.37)
Effectiveness 1.87 (1.49) 2.03 (1.29) 3.69 (0.66) 3.74 (0.60) 1.24 (1.36) 1.59 (1.33) 1.36 (1.34) 2.05 (1.38)
Explainability 0.86 (0.67) 0.64 (0.67) 2.88 (1.04) 2.70 (1.15) 0.50 (0.38) 0.36 (0.50) 0.50 (0.58) 0.49 (0.59)
Grammar 2.52 (0.89) 3.24 (0.77) 3.85 (0.37) 3.81 (0.42) 1.28 (0.99) 2.88 (1.07) 2.48 (1.03) 3.06 (0.83)
Groundness 1.48 (1.02) 2.71 (1.00) 3.74 (0.43) 3.87 (0.47) 1.08 (1.00) 2.69 (1.12) 1.04 (1.04) 2.57 (1.08)
Naturalness 1.68 (0.70) 3.27 (0.57) 3.76 (0.72) 3.74 (0.44) 1.76 (0.65) 2.90 (0.86) 2.29 (0.71) 2.81 (0.63)
Novelty 2.26 (1.25) 1.33 (1.19) 3.01 (1.14) 2.64 (1.22) 2.02 (1.42) 1.71 (1.26) 2.57 (1.43) 1.42 (1.30)
Proactiveness 1.12 (0.73) 1.26 (0.76) 3.47 (0.61) 3.43 (0.62) 0.51 (0.52) 0.81 (0.52) 0.69 (0.58) 0.69 (0.58)
Recoverability 2.33 (1.32) 2.22 (1.36) 3.91 (0.46) 3.85 (0.55) 1.22 (1.15) 2.74 (0.96) 1.11 (1.14) 1.20 (1.39)
Semantic Rel. 1.33 (1.20) 1.56 (1.28) 2.39 (1.32) 2.90 (1.06) 0.52 (0.88) 0.73 (1.07) 1.33 (1.22) 1.53 (1.36)
4.1. Experimental Setup
Dataset. Following previous work (Wang et al., 2023a, b; Zhu et al., 2024b, a), we select ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019) as the benchmark datasets. ReDial primarily focuses on movie recommendations, and OpenDialKG encompasses a broader range of domains, including movies, books, sports, and music. Table 1 shows a summary of the two datasets.
CRS Model. Following previous work (Wang et al., 2023a, b; Zhu et al., 2024b, a), we present a comparative evaluation and analysis towards four CRSs, including BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), UniCRS (Wang et al., 2022b) and CHATCRS (Wang et al., 2023a). For BARCOR, KBRD, and UniCRS, we use the pre-trained models provided by Wang et al (Wang et al., 2023a). For CHATCRS, we implemented several modifications with the GPT-4o-mini-0718 as the LLM model.
User Simulator. Building on the work of Wang et al. (Wang et al., 2023a), we use GPT-4o-mini-0718 to simulate user interactions with the CRS in dialogues. To ensure sufficient user-system interactions, we modified the interaction rules and prompts to guarantee that each dialogue session between the user simulator and CRS includes at least three and no more than five rounds of interaction. More specifically, before each session begins, we input the ground-truth item from a real user into the simulator, instructing it to describe relevant attributes during the dialogue without directly revealing the ground-truth item. In each round of dialogue, the system generates a response and a list of recommended items. The session ends if the system identifies the ground-truth item and more than three rounds of dialogue have occurred. Alternatively, the session also ends if more than five rounds of dialogue have occurred and the system has not identified the ground-truth item in any of the previous rounds.
Evaluation Process. After obtaining dialogue logs through user simulation, we post-processed the logs by consolidating the items recommended by CRS in each round into a session-level set (for the sake of convention, we refer to the set as the “recommendation list” in the paper). The left side of Figure 2 shows an example of a processed log, where all items recommended by the system during the dialogue are combined into the recommendation list. For each CRS, we obtained 3,341 records on the ReDial evaluation set and 1,187 records on the OpenDialKG evaluation set.
We then utilized LLMs to evaluate the dialogue sessions in the logs, based on the 12 factors listed in Section 3. For each factor, we have described its definition, scoring criteria, and scoring steps in the instruction. For each factor, a score between 0 and 4 was obtained. For details on the scoring, refer to Appendix A. We selected three LLMs as evaluators, namely GPT-4o-mini-0718, LLaMa3-Chat-8B, and GLM-4-Air.
4.2. Experimental Result
Table 23, and 4 show the evaluation results of four systems on two benchmark datasets, using GPT-4o-mini-0718, LLaMa3-8B, and GLM-4-Air as evaluators.
From Tables 2, 3, and 4, we can reach the following findings regarding the rankings of systems:
  1. (1)
    Across both the ReDial and OpenDialKG datasets, the three Evaluators consistently give the highest scores to CHATCRS on most metrics. Both LLaMa and GLM give CHATCRS the highest scores on all metrics, while GPT gives CHATCRS the highest scores on all factors except for Diversity.
  2. (2)
    Across both the ReDial and OpenDialKG datasets, the three evaluators consistently give BARCOR the second-highest scores on most metrics. LLaMa gives BARCOR the second-highest score on all metrics except for Novelty on ReDial; on OpenDialKG, it gives BARCOR the second-highest score on all metrics except for Diversity and Semantic Relevance. According to the scores from GPT and GLM, BARCOR ranks second in the scores for at least eight factors across both datasets.
  3. (3)
    For the systems ranked third and fourth in each factor, there is a significant amount of inconsistency in the results given by different evaluators across different datasets. For example, in most cases, both GPT and LLaMa gave UniCRS the third-highest score and KBRD the lowest score. However, this pattern does not apply to GLM. GLM gave KBRD the third-highest score in Coherence and Groundness, while giving UniCRS the lowest score. In Novelty, GLM ranked the four systems on the OpenDialKG dataset as CHATCRS ¿ UniCRS ¿ BARCOR ¿ KBRD, but on the ReDial dataset, the ranking was CHATCRS ¿ KBRD ¿ UniCRS ¿ BARCOR.
Further observation of the score rankings given by the three evaluators for the four systems, we reach the following findings reagarding the consistency (or inconsistency) of the system score rankings among the three evaluators:
  1. (1)
    In most cases, the score rankings given by GPT and LLaMa are consistent. The system score rankings for Appropriateness, Coherence, Effectiveness, Explainability, Naturalness, and Proactiveness given by GPT and LLaMa are consistent across the two datasets. For the factors of Grammar, Groundness, and Semantic Relevance, GPT and LLaMa also achieve consistent rankings for the systems on one dataset.
  2. (2)
    The score rankings given by GLM are not as consistent with those of GPT or LLaMa. GLM is only consistent with GPT and LLaMa on both datasets in terms of the score rankings for the Appropriateness factor on the OpenDialKG dataset. For the factors of Coherence, Effectiveness, Explainability, Naturalness, and Proactiveness, where GPT and LLaMa have reached consensus on the rankings across both datasets, GLM can only achieve the same conclusions as GPT or LLaMa on the ReDial dataset.
From the perspective of the datasets, we observed that GPT and LLaMa tend to provide more consistent rankings on OpenDialKG, while GPT and GLM show greater consistency on ReDial.
5. Conclusion
In this paper, we have proposed a user-centric evaluation framework based on LLMs for accurately and comprehensively evaluating conversational recommender systems. In this framework, we leverage LLM-based user simulators to interact with CRSs to obtain the interaction dialogues, and then used LLM as evaluators to assign scores to the 12 evaluation metrics of four dimensions based on the interaction dialogues. We further conduct experiments with the framework on four CRSs and two benchmark datasets to demonstrate its performance.
5.1. Main Findings
We found that the three evaluators generally awarded the highest score to CHATCRS, the second-highest to BARCOR, while the rankings for KBRD and UniCRS varied depending on the factors and datasets. Additionally, we found that in most cases, the score rankings from GPT and LLaMa were consistent; however, the rankings from GLM were less consistent with those of GPT and LLaMa.
We believe there are two possible reasons for the inconsistencies in the LLMs’ scores. First, each model learns different parameterized knowledge during training due to variations in the training corpora, which can affect its judgments. Second, certain factors, such as diversity and novelty, are more subjective in nature compared to others, such as groundedness or grammar. Such subjectivity can also influence the scores given by the LLMs, as reflected in the system rankings and the standard deviation of each system’s scores.
5.2. Limitations and Future Directions
As a pioneering work in evaluating conversational recommendation systems based on Large Language Models, this study still has several limitations that need to be addressed. Firstly, we have not examined the alignment between LLM evaluators’ scores and human ratings. Future research should incorporate user studies and human feedback to bridge this gap. Secondly, we have not yet addressed the gap in CRS evaluation where recommendation and conversational tasks are treated separately. These tasks need to be integrated into a unified approach. Additionally, previous work on search system evaluation has shown that users may be influenced by cognitive biases during their interactions with systems (Chen et al., 2023a, 2024). Future research needs to further integrate psychological theories to make the evaluation results closer to real user feedback.
Appendix A Evaluation Instructions
A.1. Instruction Template
Figure 3. The instruction templates we provide to the large models.
Figure 3 shows the instruction templates provided to the LLM evaluators. The content of FACTOR_DEFINITION, EVALUATION_STANDARD, EVALUATION_STEP for each factor, and an example of CONVERSATION_LOG can be found below.
A.2. Evaluation Guidelines
Figure 4567891011121314, and 15 show the evaluation guidelines of each factor we provided to the LLM evaluators, respectively.
Figure 4. Evaluation guideline of appropriateness
Figure 5. Evaluation guideline of coherence
Figure 6. Evaluation guideline of diversity
Figure 7. Evaluation guideline of effectiveness
Figure 8. Evaluation guideline of explainability
Figure 9. Evaluation guideline of grammatical correctnes
Figure 10. Evaluation guideline of groundness
Figure 11. Evaluation guideline of naturalness
Figure 12. Evaluation guideline of novelty
Figure 13. Evaluation guideline of proactiveness
Figure 14. Evaluation guideline of recoverability
Figure 15. Evaluation guideline of semantic relevance
Figure 16. An example of the conversation log after processing
A.3. Example of Conversation Log
Figure 16 shows an example of the conversation log after processing.